# Synchronized motion of gold nanoparticles in an optothermal trap

### v2

@@ 2.3 Numerical estimation of forces
Figure 4: Computation of Forces. a) Optical potential around the anchor AuNP calculated by integrating the optical scattering force. b) Temperature distribution around the heated anchor nanoparticle. c) The computed Optical, Opto-thermoelectric, and Thermo-osmotic slip induced drag forces on a particle near the anchor particle at various distances. The sum of all these forces along the x direction is also plotted to show the attractive behaviour at large distance.
@@
In this section, we study numerically various forces present in our system which could give rise to the observed dynamics of particles. We focus on the observation of the confinement at a radial position. The major forces present in the system are the optical forces due to the laser beam, heating-induced opto-thermoelectric force, and thermo-osmotic slip flow-induced drag force. In the presence of the laser beam, the diffusing nanoparticles experience a force due to momentum exchange with the optical field. The resultant optical force is generally decomposed into scattering and gradient force, where the gradient force is responsible for the stable trapping of particles at the beam centre. In our study, the AuNPs used in the solution have a large scattering cross section and experience a large scattering force. Consequently, they are unable to access the high-intensity region of the beam. We show the optical potential profile around the anchor particle, which is shown in Figure 4a. The potential is calculated at 1 mW of illumination power with a beam waist of around 1 micron. The optical potential is positive at all places, indicating that the in-plane force on any diffusing particle is always repulsive, pointing away from the anchor particle. The optical potential is calculated using the generalized multiparticle mie theory-based calculation as described in the methods. The potential profile is not circularly symmetric as the excitation laser is x-polarised. This gives a large force along the y-direction as compared to the x-direction.
@@
We also simulate the heating of the anchor particle by calculating the absorption coefficient of the gold nanoparticle using the Mie theory as shown in Figure 4b. Our simulated temperature rise matches well with the experimental measurement (See Supporting Information 3) of the heating of anchor nanoparticles. We can then calculate the heating of the surrounding substrate and fluid, allowing us, in turn, to calculate the fluid flow around the particle. This fluid flow is classified into two parts. The first part is convection, caused by changes in the density of the heated fluid. The second part is thermo-osmotic slip flow, resulting from differences in the free energies of the fluid across a temperature gradient parallel to the substrate. Since thermo-osmotic slip flow is the dominant contribution, we will focus on that. We modelled this thermo-osmotic slip flow using a commercial FEM-based solver using the expression from the literature $v_{||}=\chi\frac{\nabla_{||}T}{T}$, where $\chi$ is the thermo-osmotic coefficient, $T$ the temperature and $\nabla_{||}T$ is the temperature gradient parallel to the surface. We then calculate the drag force on the particle from this fluid flow using the stokes drag formula $F_{slip}=6\pi\eta Rv_{||}$, where $\eta$ is the viscosity of the fluid and $R$ is the radius of the particle.
@@
We can also calculate the opto-thermoelectric (OTE) force using the temperature profile around the particle. OTE force has been shown to be very prominent in optothermal trapping studies in the presence of surfactant. The OTE force arises in systems with surfactants and local heating, as shown in the literature. The working principle of this force is the separation of charges of the cationic surfactant in the solution. The surfactant (CTAC) molecule dissociates into the chloride ion and the CTA+ cation in the solution. For concentrations exceeding the critical micellar concentration of the surfactant, the CTA+ ions self-assemble into micelles. The CTA+ ions are also adsorbed on all particle surfaces, rendering them positively charged. In the presence of the temperature distribution, the macro-cations (micelles) and the anions start moving down the temperature gradient owing to their intrinsic thermophoresis. However, the drift speed of migration of the macro-cations is larger than the speed of anions as the anions are smaller and more diffusive. This difference in the thermo-diffusivities results in a separation of charges in the steady state. This separation of charges creates an electric field which drives the positively charged particles towards the heat centre. This mechanism can provide a large force for stable trapping in many cases, but it is not significant in our case, as can be seen in Figure 4c. The OTE force is directly proportional to the product of the temperature gradient and temperature. The temperature field decays as $\frac{1}{r}$ and so the temperature gradient decays as $\frac{1}{r^{2}}$ and thus the opto-thermoelectric force scales as $\frac{1}{r^{3}}$. Thus, the OTE force reduces in magnitude significantly compared to the other forces in the separation range of interest.
@@
We show a combination of all computed forces in Figure 4c, which shows an equilibrium position near 0.6 microns. We note that since the contribution of the OTE force is negligible, the force profile can be almost identical for the case without the surfactant, which has starkly different dynamics. We conclude this section by mentioning that while simulations show the force profile to explain the observations, they are incomplete in explaining the repulsion and synchronised motion between particles.
2.4 Emergence of spontaneous rotation
@@
Figure 5: Spontaneous rotation of trapped particles. a) Time series of particles showing the emergence of spontaneous rotation in trapped particles. b) Trajectories of particles.
@@
In this section, we experimentally demonstrate that spontaneous rotation can emerge in the trapped particle assemblies. Figure 5a shows the time series of the particles corresponding to the supplementary video 4. The particles are trapped at radial positions, as described previously. The three-particle system, as shown in figure 5a, diffuses and gets into a particular configuration due to thermal noise. After achieving a particular configuration, the particle system starts rotating suddenly. The particle system rotates for two rotations in about 5 seconds. Then, the particles rearrange again due to fluctuations, and the rotation ceases. The position distribution in Figure 5b shows that the particles cover a circular trajectory.
3 Methods
3.1 Materials
Milli Q water was used to prepare the sample. Cetyltrimethylammonium chloride (CTAC) was obtained from Sigma-Aldrich in solution (25 wtin water). Gold nanoparticles with a 400 nm diameter, stabilised suspension in citrate buffer, were purchased from Sigma-Aldrich. We have characterised the purchased particles through TEM images and UV-Vis spectra, which are shown in supplementary information. Two kinds of suspensions were created with gold nanoparticles. One in water and the other in ethanol. 1 mL of nanoparticle suspension was centrifuged, and the supernatant was removed and redispersed in 2 mL of ethanol for drop-casting on the surface. Similarly, 1 mL of nanoparticle suspension was centrifuged and redispersed in 1 mL of MilliQ water. Two 50 µL solutions were prepared in micro-centrifuge tubes using MilliQ water to prepare nanoparticle suspension in surfactant solution. One with 5 µL of previously prepared aqueous gold NP suspension added to 45 µL water. Another tube had 10 µL of freshly prepared 50 mM CTAC concentration solution added to 40 µL of water. These 50 µL solutions were mixed to give the final solution, which was used for trapping. Goldseal Cover Glass from Ted Pella, Inc. were used for experiments after cleaning with acetone. A double-sided adhesive spacer from Grace Bio-Laboratories (SecureSeal imaging spacer, 120 µm thickness) was used to build a microfluidic chamber as follows. The spacer is pasted on the cover glass. 2 µL of the alcoholic gold suspension was drop-casted on the glass coverslip and allowed to dry by evaporation for 20 minutes. This gives gold NPs which are firmly attached to the surface. After 20 minutes, a 10 µL solution of the surfactant mixed with aqueous suspension is poured into the spacer, and the chamber is sealed with another cover glass from the top.
3.2 Optical Measurements
The experiments were performed on a custom-built dual-channel microscope. The imaging was performed in brightfield mode, and trapping was performed in upright geometry. A 100$\times$, 0.95 NA objective lens (Olympus MPlanApo N) focused the laser from the top of the AuNP anchored on the cover glass. The same objective also illuminates the sample plane with white light from the top. Another 100$\times$, 1.49 NA objective lens (Olympus UApo N, oil immersion) was used to collect all the transmitted light and the images were captured using a fast camera with a frame rate of 500. The laser wavelength is filtered from the collected light before reaching the camera using a notch filter and a long pass filter. The setup was equipped with a three-dimensional piezoelectric translation stage for high-precision control of the sample.
3.3 Particle Tracking
A fast camera recorded the particle trajectories at a frame rate of 500 frames per second. The images were tracked using the Trackmate plugin of the FiJi distribution of ImageJ73, 74. The images shown were taken using brightfield illumination, and the gold nanoparticles are shown as dark objects on a bright background. These images were then inverted for tracking using TrackMate plugin. The images extracted from the tracked processed video show particles as white objects on a black background.
3.4 Numerical Calculations
We calculated the scattering forces between particles using a Python-based generalized multiparticle mie theory package, MiePy. The forces were then integrated to get the potential profile (see Supporting Information 5). The absorption cross-section of the nanoparticle was calculated using Mie theory, which was used to calculate the temperature rise of the nanoparticle. The temperature distribution around the nanoparticle was then calculated by solving the Heat-diffusion equation by convolving the heat source density with the Green’s function. Using the temperature rise of the nanoparticle, we conducted numerical simulations using a finite element method (FEM) solver to calculate the thermo-osmotic slip flow-induced force (see Supporting Information 6). The OTE force is similarly computed using the temperature profile as shown in Supporting Information 7.
@@ 4 Conclusion
This study investigates the dynamics of nanoparticles in anchor particle-based optothermal traps, focusing on the behaviour of gold nanoparticles in surfactant solutions. We observed that particles are trapped some distance from the centre and exhibit synchronized rotational diffusion and spontaneous rotation around a heated anchor particle. The dynamics suggest the presence of some repulsive interparticle force, which helps synchronise the particles’ motion even with interparticle separation of around 1 micron. These unique dynamics open possibilities for precisely tuning particle interactions, which could be highly valuable for applications in nanotechnology, such as assembling nanoparticle arrays or manipulating particles in nanoscale systems.
The presence of surfactants, particularly in the form of CTAC solutions, plays a critical role in influencing the trapping and motion of nanoparticles. Our results suggest that surfactants help confine particles to specific radii around the anchor particle, leading to organized rotational diffusion. This behaviour was absent when either the anchor particle or the surfactant was removed, underscoring its importance in achieving controlled trapping. Overall, our study aims to contribute to a deeper understanding of particle dynamics in complex trapping environments, offering insights that could be used to design new nanoparticle assembly techniques. These findings hold potential for advancing applications in materials science, nanofabrication, and optoelectronics, where controlled particle interaction and organization are important. {acknowledgement} The authors thank Dr Adarsh Vasista, Dr Sunny Tiwari and Dr Diptabrata Paul for their valuable discussions on this project. We acknowledge using common experimental facilities in the Chemistry Department at IISER Pune. AS acknowledges the Ministry of Education, Government of India, for the Prime Minister Research Fellowship. This work was partially funded by AOARD (grant number FA2386-23-1-4054) and the Swarnajayanti fellowship grant (DST/SJF/PSA-02/2017-18) to G.V.P.K.
@@
Comparison of particle distribution and dynamics with and without CTAC. Two particle synchronized rotational diffusion. Measurement of temperature using liquid crystal phase transition. Computation of temperature rise and distribution. Computation of optical potential. Computation of thermo-osmotic slip flow and force. Computation of opto-thermoelectric force. (pdf)
Supplementary Video 1: Experimental video showing trapping and synchronization of gold particles. Supplementary Video 2: Comparision of trapping dynamics in various conditions. a) without anchor particle and with CTAC, b) without CTAC and with anchor particle, c) with anchor particle and 5 mM CTAC. Supplementary Video 3: Experimental video with simultaneous angular position graph showing synchronised motion. Supplementary Video 4: Experiment video showing spontaneous rotation in trapped particle assembly.

### v3

@@ 3.3 Numerical Estimation of Forces
Figure 4: Computation of Forces. a) Optical potential around the anchor AuNP calculated by integrating the optical scattering force. b) Temperature distribution around the heated anchor nanoparticle. c) The computed Optical, Opto-thermoelectric, and Thermo-osmotic slip induced drag forces on a particle near the anchor particle at various distances. The sum of all these forces along the x direction is also plotted to show the attractive behaviour at a large distance.
@@
In this section, we study numerically various forces present in our system which could give rise to the observed dynamics of particles. We focus on the observation of the confinement at a radial position. The major forces present in the system are the optical forces due to the laser beam, heating-induced opto-thermoelectric force, and thermo-osmotic slip flow-induced drag force. In the presence of the laser beam, the diffusing nanoparticles experience a force due to momentum exchange with the optical field. The resultant optical force is generally decomposed into scattering and gradient force, where the gradient force is responsible for the stable trapping of particles at the beam centre. In our study, the AuNPs used in the solution have a large scattering cross section and experience a large scattering force. Consequently, they are unable to access the high-intensity region of the beam. We show the optical potential profile around the anchor particle, shown in Figure 4a. The potential is calculated at one mW of illumination power with a beam waist of around 1 $\mu m$. The optical potential is positive at all places, indicating that the in-plane force on any diffusing particle is always repulsive, pointing away from the anchor particle. The optical potential is calculated using the generalized multiparticle Mie theory-based calculation described in the methods. The potential profile is not circularly symmetric as the excitation laser is x-polarised. This gives a large force along the y-direction as compared to the x-direction.
@@
We also simulate the heating of the anchor particle by calculating the absorption coefficient of the gold nanoparticle using the Mie theory as shown in Figure 4b. Our simulated temperature rise matches well with the experimental measurement (See Supporting Information S2 and S3) of the heating of anchor nanoparticles. We can then calculate the heating of the surrounding substrate and fluid, allowing us, in turn, to calculate the fluid flow around the particle. This fluid flow is classified into two parts. The first part is convection, caused by changes in the density of the heated fluid. The second part is thermo-osmotic slip flow, resulting from differences in the free energies of the fluid across a temperature gradient parallel to the substrate. We focus on thermo-osmotic slip flow as it is the dominant contribution. We modelled this thermo-osmotic slip flow using a commercial FEM-based solver using the expression from the literature $v_{||}=\chi\frac{\nabla_{||}T}{T}$, where $\chi$ is the thermo-osmotic slip coefficient, $T$ the temperature and $\nabla_{||}T$ is the temperature gradient parallel to the surface. We then calculate the hydrodynamic drag force on the particle from this fluid flow using the Stokes drag formula $F_{slip}=6\pi\eta Rv_{||}$, where $\eta$ is the viscosity of the fluid, and $R$ is the radius of the particle.
@@
We can also calculate the opto-thermoelectric (OTE) force using the temperature profile around the particle. OTE force is very prominent in optothermal trapping studies in the presence of the surfactant. The OTE force arises in systems with surfactants and local heating, as shown in the literature. The working principle of this force is the separation of charges of the cationic surfactant in the solution. The surfactant (CTAC) molecule dissociates into the chloride ion and the CTA+ cation in the solution. For concentrations exceeding the critical micellar concentration of the surfactant, the CTA+ ions self-assemble into micelles. The CTA+ ions are also adsorbed on all particle surfaces, rendering them positively charged. In the presence of the temperature distribution, the macro-cations (micelles) and the anions start moving down the temperature gradient owing to their intrinsic thermophoresis. However, the drift speed of migration of the macro-cations is larger than that of anions as the anions are smaller and more diffusive. This difference in the thermo-diffusivities results in a separation of charges in the steady state. This separation of charges creates an electric field which drives the positively charged particles towards the heat centre. See supporting information S5 for further details. This mechanism can provide a large force for stable trapping in many cases, but it is not significant in our case, as seen in Figure 4c. The OTE force is directly proportional to the product of the temperature gradient and temperature. The temperature field decays as $\frac{1}{r}$ and so the temperature gradient decays as $\frac{1}{r^{2}}$ and thus the opto-thermoelectric force scales as $\frac{1}{r^{3}}$. Thus, the OTE force reduces in magnitude significantly compared to the other forces in the separation range of interest.
@@
We combine all computed forces in Figure 4c, which shows an equilibrium position near 0.6 $\mu m$. Since the contribution of the OTE force is negligible, the force profile would be almost identical for the case without the surfactant, which has starkly different dynamics. Thus, we can infer that the surfactant facilitates some additional interactions which are not yet understood in the literature.
We would like to mention that due to the presence of the concentration gradient of surfactant micelles, there could also be a depletion interaction. However, this depletion interaction has been shown to be weaker than the OTE force in surfactant-assisted Optothermal traps. 49 Thus, we have not considered its contribution to our analysis.
To conclude, we would like to mention that the forces responsible for the synchronized motion are poorly understood. Hydrodynamic synchronization in micron-sized, dielectric, optically driven colloids has been shown in various contexts.81, 82, 83 We believe such hydrodynamic synchronization could be the possible reason in our experiments as well. However, the AuNPs do not show any hydrodynamic correlation in the absence of the surfactant. Owing to a multitude of factors, such as the presence of laser scattering from the anchor particle, the surfactant micelle concentration gradient and the trapping close to the surface, where micelles are adsorbed, the problem here is more complex, and possibly a mixture of multiple effects to understand it. Since the trapped particles also experience laser illumination since the trapping radii are of the order of the beam waist, there would also be finite heating of the trapped AuNPs. This heating would modify the thermo-osmotic flows, further complicating the analysis. Additionally, the system is highly charged due to the presence of CTA+ micelles and corresponding Chlorine anions. This high-charge environment could facilitate electrostatic repulsion-like force mediated by surfactant micelles.
In the following subsection, we discuss the experimental observation of rotation of the trapped assemblies and the dependence of rotation on experimental parameters.
3.4 Rotation of Trapped Assembly
@@
Figure 5: Rotation of the trapped particle assembly. a) Time series of particles showing the emergence of spontaneous rotation in trapped particles. b) Trajectories of particles.
@@
The nanoparticle assemblies that are formed tend to rotate around the anchor particle. This rotation can be seen in all supplementary videos, and the rotation direction is stochastic. In this section, we experimentally demonstrate that faster spontaneous rotation can emerge in the system. Figure 5a shows the time series of the particles corresponding to the supplementary video 6. The particles are trapped at radial positions, as described previously. As shown in Figure 5a, the three-particle system diffuses and gets into a particular configuration due to thermal noise. After achieving a particular configuration, the particle system starts rotating suddenly. The particle system rotates for two rotations in about 5 seconds. Then, the particles rearrange again due to fluctuations, and the rotation ceases. The position distribution in Figure 5b shows that the particles cover a circular trajectory.
Recently, Louis et al. have also shown trapping and correlated motion of 200 nm AuNPs using a 1064 nm laser beam.84 In particular, they show correlated rotative oscillations for three trapped particles. In their case, stable optical trapping is possible because of the particle-laser combination they experimented with, viz. 200 nm diameter AuNP with a 1064 nm laser beam. This is similar to what Tao et al. used to show indirect optical binding.41 The particle separation in their experiments is of the order of a wavelength in the medium. In contrast, the particles in our experiments show synchronised motion at even separation of the order of 4 times the laser wavelength in the medium, as shown in Figure 3. We also do not observe any stable conformation of the particles. All assemblies are formed stochastically based on the direction of the particle approach.
We observe this rotation using linearly polarised light, which carries no spin angular momentum (SAM). We further probed the effect of light with SAM on the assembly’s rotation direction, as has been observed for individual particle plasmonic trapping in literature.20 We found that the angular momentum of the light does not influence the rotation direction of the assembly. A representative experiment is shown in supplementary video 7 and described in Figure 6. The particles are trapped as described in the subsection ‘Anchor particle driven trapping’ but using a circularly polarised laser beam. We use a quarter wave plate to make the incident light circularly polarised.
Figure 6: Trapping with circularly polarised light. a) shows the trajectory of one trapped particle from the supplementary video 7 for 40 seconds. The angular position of the particle is shown in b). The red dotted lines mark the time points when a new particle joins the assembly. Using a fixed circular polarisation does not tend to induce rotation in a preferential direction.
The trajectory of one trapped particle is shown in Figure 6a instead of all particles in the assembly for clarity. The corresponding angular trajectory is shown in 6b for 40 seconds. The assembly rotates in the counter-clockwise direction for more than four rotations in 15 seconds, then the rotation stagnates. More particles start joining the assembly, and the rotation direction slowly reverses. The multi-particle assembly performs roughly one clockwise rotation in the next 25 seconds. The red dotted lines in Figure 6b are the time points when a new particle joins the assembly. Thus, the rotation direction of the assembly does not depend explicitly on the laser polarisation. Therefore, the observed rotation direction is stochastic and depends on the specific conformation of the assembly.
@@ 4 Conclusion
This study explores the dynamics of nanoparticles in anchor particle-based optothermal traps, focusing on the behaviour of 400 nm AuNPs in CTAC surfactant solution. We observed that nanoparticles are not trapped at the beam focus. Instead, we confined them at micrometer-scale distances from the heated anchor particle, where they exhibit synchronised rotational diffusion and spontaneous rotation. The presence of a repulsive interparticle force appears to regulate their motion, enabling synchronisation even at separations of several micrometres. The results highlight the critical role of CTAC in shaping nanoparticle interactions in an optothermal trap. Our results indicate that radial confinement and coordinated motion are absent in the absence of either the anchor particle or surfactant. This suggests that additional forces beyond known optical and thermophoretic mechanisms contribute to the observed trapping behaviour. Further studies are required to understand the precise mechanism of the synchronisation.
By studying these novel trapping dynamics, this work aims to expand our understanding of optothermal trapping in surfactant environments. Our experiments demonstrate that nanoparticle confinement and synchronised motion depend on key parameters such as surfactant concentration, anchor particle size, and laser power. We observed that reducing the surfactant concentration below the critical micellar concentration led to closer trapping near the anchor particle and weaker angular correlations. Varying the size of the anchor particle slightly altered the trapping radius, while increasing laser power steadily increased the mean trapping radius. We also observe the rotation of trapped particle assemblies. Experiments with circularly polarised light confirmed that the spontaneous rotation of particle assemblies is not driven by the optical spin angular momentum but instead emerges from the trapped assemblies’ specific conformations.
These results collectively highlight the complex interplay of optical, thermal, and surfactant-mediated forces, paving the way for new approaches and applications in materials science, nanofabrication, and optoelectronics, where controlled particle interaction and organisation are essential. {acknowledgement} The authors thank Dr Adarsh Vasista, Dr Sunny Tiwari and Dr Diptabrata Paul for their valuable discussions on this project. We acknowledge using common experimental facilities in the Chemistry Department at IISER Pune. AS acknowledges the Ministry of Education, Government of India, for the Prime Minister Research Fellowship. This work was partially funded by AOARD (grant number FA2386-23-1-4054) and the Swarnajayanti fellowship grant (DST/SJF/PSA-02/2017-18) to G.V.P.K.
@@
Computation of optical potential, Measurement of temperature using liquid crystal phase transition, Computation of temperature rise and distribution, Computation of thermo-osmotic slip flow and force, Computation of opto-thermoelectric force, Optical binding comparison, Comparison of particle distribution and dynamics with and without CTAC, Variation of anchor particle size, Variation of surfactant concentration, Laser power dependence on trapping behaviour, Two particles synchronised rotational diffusion (pdf)
Supplementary Video 1: Experiment video showing trapping and synchronisation of gold particles. Supplementary Video 2: Comparison of trapping dynamics in various conditions. a) without anchor particle and with CTAC, b) without CTAC and with anchor particle, c) with anchor particle and five mM CTAC. Supplementary Video 3: Experiment video comparing the trapping with variation of anchor particle size. Supplementary Video 4: Experiment video comparing the trapping with variation of surfactant concentration. Supplementary Video 5: Experiment video with simultaneous angular position graph showing synchronised motion. Supplementary Video 6: Experiment video showing spontaneous rotation in trapped particle assembly. Supplementary Video 7: Experiment video showing the effect of circularly polarised light in rotation of trapped particle assembly.